# Preference-Guided Reinforcement Learning for Efficient Exploration

### v1

@@ VII Experimental Results
This section compares LOPE’s performance with several state-of-the-art baseline methods on extensive benchmark control tasks. These average returns were calculated over ten separate runs with different random seeds, and the shaded error bars represented the standard errors. We evaluate the agent performance quantitatively by measuring the average return computed by ground truth rewards of environments for fairness.
VII-A Grid World
[…]
Figure 3: (a) Learning curves of success rate in the grid-world maze; (b) Changing trend of MMD distance.
As shown in Fig. 3a, LOPE outperforms the other baseline methods in this grid-world task. Specifically, LOPE can learn faster and achieve a higher success rate. We noted that for a long time, the baseline methods, such as PPO and SIL, could not even learn a sub-optimal policy that picks up the key and then opens the door. PPO+EXP learned faster than PPO and SIL because this algorithm could explore the environment more efficiently and find the treasure. Interestingly, LOPE learned the optimal policy most quickly. LOPE exploited preference-labeled experiences and quickly learned to open the door and further explore the broader state space. This increased the chance of obtaining the treasure reward and helped the agent learn the optimal policy during training. This instance indicates that LOPE can drive deep exploration and improve the learning efficiency of the policy with some human feedback.
@@ VII-A Grid World
Fig. 3b shows the changing trend of the MMD distance between $\pi$ and $\pi_{b}$ during the training process of LOPE. Except for the later stages of the training process, a considerable MMD distance was maintained between $\pi$ and $\pi_{b}$. Intuitively, the policy $\pi_{b}$ implied by the preference data is superior to the current policy $\pi$. Therefore, this result demonstrates the reasonableness of Assumption 1. In this manner, updating the policy parameters towards $\pi_{b}$ guarantees a higher performance improvement bound. We further depicted the actual and human-annotated trajectories in two state visitation graphs. The results are presented in Fig. 4. It implies that the agent can imitate the preference-labeled trajectories perfectly. Therefore, our method can exploit human preferences efficiently and learn the desired policy.
VII-B Locomotion Tasks from MuJoCo
For easier access to preference-labeled trajectories, and notice that SparseHalfCheetah and SparseHopper can only move along the $x$-axis, we set up three different nodes roughly evenly within the range of motion of the agent to determine preference-labeled trajectories. Trajectories that passed through these nodes more “gracefully” and “faster” were preferred. Specifically, the term “graceful” is used to describe that an agent can travel a greater distance at each step with less cost, that is, minimizing control costs. The term “fast” indicates that a trajectory can approach the target in fewer steps, that is, minimizing the number of steps.
@@ VII-B Locomotion Tasks from MuJoCo
As shown in Fig. 5, by selecting the preferred trajectories through the set nodes and considering them as guidance, the LOPE agent can reproduce similar behaviors quickly and gather the survival bonus in both sparse reward tasks. On the other hand, although the PEBBLE and GASIL agents exploited previous good experiences similar to our method, these methods found the optimal policy to obtain the sparse reward more slowly. They even failed to learn the optimal policy in these hard-exploration tasks. These approaches do not exploit the information in preferences, such as the agent’s coordinate position in the environment, and create remarkable information bottlenecks when utilizing previous good experiences. This information bottleneck prevents the agent from directly learning from these demonstrations. Furthermore, the PPO agent cannot perform satisfactorily on the SparseHalfCheetah and SparseHopper tasks. They depend on the action- and parameter-space noise sampled from the Gaussian distribution, which is inadequate and inefficient for hard-exploration tasks.
VII-C Hard Exploration MuJoCo Maze
We set three different distance nodes and consider trajectories passing farther distance nodes as preference-labeled trajectories with higher priority in the ant-maze task. These nodes are the two corners and the target point. Trajectories that passed through these nodes more “gracefully” and “faster” were preferred. Fig. 6a presents the result.
Figure 6: (a) Learning curves of success rate on the Ant-Maze task; (b) Changing trend of MMD distance.
Compared with the other baseline methods, LOPE can learn faster and achieve higher convergence values of success rate. LOPE has a solid theoretical foundation that provides a guarantee for performance improvement. Moreover, PEBBLE never received a positive reward on this challenging exploration task and did not benefit from the unsupervised pre-training and learned reward function consistent with preferences. Meanwhile, we found that SIL performed much worse than LOPE during training. This algorithm cannot encounter a positive reward due to random exploration. This made it impossible to learn a good policy because there was no previous good demonstration to exploit. Furthermore, NoisyNet and PPO failed to reach the goal and learn a good policy within a reasonable time frame. This result indicates that an advanced exploration strategy is essential in hard-exploration environments for RL algorithms, and LOPE can exploit preference-labeled trajectories as guidance to drive deep exploration. Hence, our approach can avoid the potential information bottleneck and fully use data information in preferences, such as the agent’s coordination position in the environment. Fig. 6b shows the changing trend of the MMD distance between $\pi$ and $\pi_{b}$ during the training process of LOPE. The changing trend is similar to that in Fig. 3b.
@@ VII-D Evaluation of Ground-Truth Reward Learning
A deeper look into experimental results is provided to verify the LOPE’s effectiveness in learning near-optimal behaviors. In this experiment, we evaluated actions learned by different agents with the ground-truth reward function, which is the default setting of OpenAI Gym [60]. We then calculated the average ground-truth reward for each agent and compared them in Fig. 7. For comparison, we also trained PPO with this dense ground-truth reward function.
According to Fig. 7, the LOPE’s average ground-truth reward dramatically increases at the beginning of training, similar to PPO’s learning trend in the default OpenAI Gym reward setting. Meanwhile, the value of the LOPE’s average ground-truth reward is improved incrementally during policy optimization, and this value gradually approximates that of PPO with the default reward function at the end of training. This result indicates that human feedback can help the agent learn near-optimal behaviors that are rewarded a high score by the default OpenAI Gym reward function.
@@ VII-E Ablation Analyses
The experimental results presented in the previous section indicate LOPE outperforms other baseline approaches on several challenging tasks. We are now interested in whether these advantages still hold when the preference setting changes. For this purpose, we designed an ablation experiment and compared the LOPE’s performance on preferences with different qualities. Moreover, we conducted another ablation study on the algorithm structure to demonstrate the effectiveness of preference guidance. These ablation studies are based on the grid world environment.
@@
LOPE with different structures. We ran LOPE without the policy improvement step (LOPE w/o PI) and LOPE without the preference guidance step (LOPE w/o PG) and obtained their results separately. The results are presented in Fig. 8LABEL:sub@fig:ablation. LOPE w/o PG cannot find the treasure and learn the optimal policy. In contrast, LOPE w/o PI can find the treasure; however, the learning efficiency of the optimal policy is low. This result implies that preference guidance can improve the agent’s exploration efficiency while the policy improvement step guarantees the reproduction of these successful trajectories.
Figure 8: Ablation results of algorithm structure and mislabeling preferences.
Preferences with different qualities. In this experiment, preference-labeled trajectories can be mislabeled with different ratios. Specifically, we selected four different mislabeling ratios, $0,0.1,0.2,0.3$, and obtained corresponding experimental results. The performance comparison is shown in Fig. 8LABEL:sub@fig:mislabel. LOPE achieves the highest performance when the mislabeling ratio is 0. This result implies that the quality of demonstrations will significantly affect the performance of LOPE, and high-quality human preferences can facilitate policy optimization and accelerate learning to some extent.
@@ VIII Conclusion
This study develops a practical PbRL algorithm for hard-exploration tasks with long horizons and sparse rewards. Our key idea is to adjust the agent’s exploration by considering preferred trajectories as guidance, avoiding learning a reward function. LOPE involves a two-step sequential policy optimization process, including trust-region-based policy improvement and preference guidance steps. Preference guidance is reformulated as a novel trajectory-wise state marginal matching problem that minimizes a novel maximum mean discrepancy distance between the preferred trajectories and the learned policy. We further provide a theoretical analysis to characterize the performance improvement bound and evaluate the effectiveness of the proposed approach. Extensive experimental results verify the superior performance of LOPE over different competitive baseline methods on several sparse-reward benchmark tasks.
Acknowledgments
This work was supported by the National Science and Technology Major Project(2022ZD0116401), the National Natural Science Foundation of China (62141605), and Fundamental Research Funds for the Central Universities.
@@ Appendix A Proof of Lemma 1
$\underset{\tau\sim\rho_{\theta}}{\mathbb{E}}\left[\underset{\upsilon\sim\mathcal{P}}{\mathbb{E}}\left[k(\tau,\upsilon)\right]\right]=\underset{\begin{subarray}{l}s\sim d_{\theta}\\ a\sim\pi_{\theta}\end{subarray}}{\mathbb{E}}\left[\underset{\upsilon\sim\mathcal{P}}{\mathbb{E}}\left[k(\tau,\upsilon)\right]\right],$ (17)
where $d_{\theta}$ is the discounted state visitation distribution defined in Section III.
In this manner, $\underset{\begin{subarray}{c}\tau\sim\mathcal{P}_{s,a}\end{subarray}}{\mathbb{E}}\left[\mathrm{dist}(\tau,\mathcal{P})\right]$ in Eq. (17) can be viewed as an intrinsic reward and let $r_{g}(s,a)=\underset{\begin{subarray}{c}\tau\sim\mathcal{P}_{s,a}\end{subarray}}{\mathbb{E}}\left[\mathrm{dist}(\tau,\mathcal{P})\right]$. Then, we obtain
$\underset{\begin{subarray}{l}s\sim d_{\theta}\\ a\sim\pi_{\theta}\end{subarray}}{\mathbb{E}}\left[r_{g}(s,a)\right]=\underset{\tau\sim\rho_{\theta}}{\mathbb{E}}\left[\underset{\upsilon\sim\mathcal{P}}{\mathbb{E}}\left[k(\tau,\upsilon)\right]\right].$ (18)
@@ Appendix B Proof of Policy Performance Bound
Definition 1 ($\alpha$-coupled policies [51, 53]).
$(\pi,\tilde{\pi})$ are $\alpha$-coupled policies if they defines a joint probability distribution $(a,\tilde{a})|s$, such that $P(a\neq\tilde{a}|s)\leq\alpha$ for all $s$. $\pi$ and $\tilde{\pi}$ will denote the marginal distributions of $a$ and $\tilde{a}$, respectively.
@@
Lemma 3 (Adopted from [61]).
@@
Lemma 4 (Adopted from [51, 54]).
@@
The proof details can be found in [51], and we omit them here. Given two arbitrary stochastic policies $\pi_{1}$ and $\pi_{2}$, define the expected advantage of the policy $\pi_{1}$ over the policy $\pi_{2}$ at state $s$ as $A^{\pi_{1}|\pi_{2}}(s)$:
@@
Here, Eq. (23) is obtained by defining the expected return $J(\tilde{\pi})$ of $\tilde{\pi}$ over $\pi_{b}$ as $J(\pi_{b})+\mathbb{E}_{\tilde{\pi}}\left[\sum^{\infty}_{t=0}\gamma^{t}A^{\tilde{\pi}|\pi_{b}}(s_{t})\right]$. If we define the expected returns of $\tilde{\pi}$ and $\pi$ over $\pi$, then we have:
$J(\tilde{\pi})$ $=J(\pi)+\mathbb{E}_{\tau\sim\rho_{\tilde{\pi}}}\left[\sum^{\infty}_{t=0}\gamma^{t}A^{\tilde{\pi}|\pi}(s_{t})\right],$ (24)
@@
A similar result can be found in [53].
Lemma 5 (Adopted from [53]).
@@
The proof details of Lemma 5 can be found in [53], and we omit them here.
@@
Now, we will derive the performance improvement bound described in Lemma 2. The content of Lemma 2 is described below.
See 2
Proof.
Since $\beta=D_{\rm KL}^{\max}(\tilde{\pi},\pi_{b})=\max_{s}D_{\rm KL}(\tilde{\pi}(\cdot|s),\pi_{b}(\cdot|s))$, considering $D_{\rm TV}^{2}(p,q)\leq D_{\rm KL}(p,q)$ [62], we assume $\alpha=D_{\rm TV}^{\max}(\tilde{\pi},\pi_{b})$ and $\alpha$ satisfies $\alpha^{2}\leq\beta$. According to Lemma 3, $\pi_{b}$, $\tilde{\pi}$ are $\alpha$-coupled. Furthermore, it is worth mentioning that the expert policy $\pi_{b}$ satisfies Assumption 1, which means
@@
Consider introducing $J(\pi_{b})$ into $J(\tilde{\pi})$ and $J(\pi)$ and applying Lemma 4 to $\tilde{\pi}$ and $\pi_{b}$, then we have
@@
$\geq$ $-\sum^{\infty}_{t=0}\gamma^{t}\left(2\alpha(1-(1-\alpha)^{t})\max_{s,a}|A_{\pi_{b}}(s,a)-\Delta\right)$ (36)
@@
where $\epsilon_{b}=\max_{s,a}|A_{b}(s,a)|$ is the maximum absolute value of $\pi_{b}$’s advantage function $A_{b}(s,a)$. Noting $\beta=D_{\rm KL}^{\max}(\tilde{\pi},\pi_{b})$ and applying $D_{\rm TV}^{2}(p,q)\leq D_{\rm KL}(p,q)$ [62] gives
@@
Lemma 2 is proved. ∎
TABLE I: Hyper-parameter choices of LOPE for different tasks.
Hyper-parameter Grid World Ant HalfCheetah Hopper
Value Function Learning Rate $0.01$ $0.001$ $0.001$ $0.001$
Policy Learning Rate $0.000015$ $0.00009$ $0.00009$ $0.0004$
Actor Regularization None None None None
Policy Optimizer Adam Adam Adam Adam
Value Function Optimizer Adam Adam Adam Adam
Max Length per Episode $240$ $750$ $500$ $500$
Episodes per iteration $8$ $30$ $20$ $20$
Epochs per iteration $80$ $65$ $80$ $80$
Discount Factor $0.99$ $0.99$ $0.99$ $0.99$
Normalized Observations False False False False
Gradient Clipping False False False False
Initial Exploration Policy None $N(0,0.17)$ $N(0,0.15)$ $N(0,0.65)$
Clipped Epsilon $0.3$ $0.21$ $0.2$ $0.27$
See 1
Proof.
@@
Appendix C Neural Network Architectures and Hyper-Parameters
C-A LOPE Policy Architecture for Discrete Control Tasks

### v2

@@ VII Experimental Results
This section compares LOPE’s performance with several state-of-the-art baseline methods on extensive benchmark control tasks. These average returns were calculated over ten separate runs with different random seeds, and the shaded error bars represented the standard errors. We evaluate the agent’s performance quantitatively by measuring the average return, computed using ground truth rewards from environments, for fairness.
VII-A Key-Door-Treasure Domain
In the Key-Door-Treasure domain, we designate the positions of the key, door, and treasure room entrance as nodes. Trajectories containing events such as picking the key, opening the door, or entering the treasure room are considered preference trajectories. Moreover, trajectories closer to the goal with fewer steps are preferred. LOPE only maintains the top-$h$ trajectories and leverages them to guide policy optimization; we set $h=5$ in all experiments.
Figure 5: (a) Learning curves of success rate in the grid-world maze with a fixed goal; (b) Learning curves of success rate in the grid-world maze with random goals.
As shown in Figs. 5a and 5b, in both grid-world mazes with fixed and random goals, LOPE’s performance is superior or equal to other baseline methods. Specifically, LOPE can learn faster and achieve a success rate of nearly 95%. We noted that in the grid-world maze with the fixed goal, the baseline methods, such as PPO and SIL, could not even learn a sub-optimal policy that picks up the key and then opens the door for a long time. PPO+EXP learned faster than PPO and SIL because this algorithm could explore the environment more efficiently and find the treasure. Interestingly, all baseline methods obtain better performance in the random grid-world environment than in the fixed grid-world maze. The reason for this phenomenon is that in the random grid-world maze, the goal generated in the upper-right room can be reached easily. In both grid-world mazes, LOPE exploited preference-labeled experiences and quickly learned to open the door and further explore the broader state space. This increased the chance of obtaining the treasure reward and helped the agent learn the optimal policy during training.
@@ VII-A Key-Door-Treasure Domain
Figure 7: (a) Learning curves of success rate on the continuous maze tasks: (1) AntMaze-Umaze; (2) PointMaze-Medium; (3) PointMaze-Large.
We further depicted the actual and human-annotated trajectories in two state visitation graphs. The results are presented in Fig. 6. It implies that the agent can imitate the preference-labeled trajectories perfectly. Therefore, our method can exploit human preferences efficiently and learn the desired policy.
VII-B Hard-Exploration Continuous Mazes
We set three different distance nodes and consider trajectories passing farther distance nodes as preference-labeled trajectories with higher priority in the continuous maze tasks. Specifically, in the ant maze, these nodes are the two corners and the target point. In the PointMaze-Medium task, these nodes are set up in sequence as $\{(3,3),(6,4),(6,6)\}$. Finally, we select $\{(3,6),(5,6),(7,9)\}$ as check nodes in the PointMaze-Large task. Trajectories that passed through these nodes more “gracefully” and “faster” were preferred. Fig. 7a presents the result.
Compared with the other baseline methods, LOPE can learn faster and achieve higher convergence values of success rate. LOPE has a solid theoretical foundation that provides a guarantee for performance improvement. Specifically, in the AntMaze environment, as shown in Fig. 7a, only the LOPE agent can reach the target point and obtain the reward. At the same time, all other baseline methods failed to learn the policy leading to the reward. In the PointMaze tasks, as shown in Figs. 7b and 7c, the LOPE achieves a higher success rate than other baseline approaches. Notably, the SMM method used a policy-level state marginal matching objective to encourage exploration, and LOPE significantly outperformed SMM, demonstrating the effectiveness of the trajectory-wise state marginal matching objective. The results compared with PEBBLE and RUNE indicate that our approach can avoid the potential information bottleneck and fully utilize the data information in preferences.
VII-C Locomotion Tasks from MuJoCo
For easier access to preference-labeled trajectories, we note that SparseHalfCheetah and SparseHopper can only move along the $x$-axis. To determine preference-labeled trajectories, we set up three different nodes roughly evenly within the range of motion of the agent. Trajectories that passed through these nodes more “gracefully” and “faster” were preferred. Specifically, the term ”graceful” is used to describe an agent that can travel a greater distance at each step with less cost, thereby minimizing control costs. The term“fast” indicates that a trajectory can approach the target in fewer steps, thereby minimizing the number of steps.
@@ VII-C Locomotion Tasks from MuJoCo
As shown in Fig. 8, by selecting the preferred trajectories through the set nodes and considering them as guidance, the LOPE agent can reproduce similar behaviors quickly and gather the survival bonus in both sparse reward tasks. Specifically, in SparseHopper, LOPE achieves performance comparable to PPO with dense rewards, indicating that the policy can approach optimality even when guided only by preferences. Due to the complexity of high-dimensional dynamics, PPO exhibits suboptimal convergence in the HalfCheetah environment. LOPE shows robust performance but also exhibits a plateau, consistent with prior observations. We hypothesize that this is due to limited exploration coverage and propose off-policy extensions to address this in future work.
On the other hand, although the PbRL methods, PEBBLE and RUNE, exploited human preferences similar to our method, these methods learned the optimal policy more slowly. These approaches do not leverage the information in preferences, such as the agent’s coordinate position in the environment, and create significant information bottlenecks when utilizing previous successful experiences. This information bottleneck prevents the agent from directly learning from these demonstrations. The SMM agent performs unsatisfactorily on the SparseHalfCheetah and SparseHopper tasks, further demonstrating the effectiveness of the trajectory-wise state marginal matching objective. Other baseline methods depend on the action- or parameter-space noise sampled from the Gaussian distribution, which is inadequate and inefficient for hard-exploration tasks.
@@ VII-D Evaluation of Ground-Truth Reward Learning
A deeper examination of the experimental results is provided to verify the effectiveness of LOPE in learning near-optimal behaviors. In this experiment, we evaluated actions learned by different agents with the ground-truth reward function, which is the default setting of OpenAI Gym [1606.01540]. We then calculated the average ground-truth reward for each agent and compared them in Fig. 9. For comparison, we also trained PPO with this dense ground-truth reward function.
Figure 10: (a) Agent trained by LOPE; (b) Agent trained by PPO with the ground-truth reward function.
Figure 11: (a) Ablation results of algorithm structure; (2) Performance comparison under mislabeling preferences; (3) Ablation results of exchanging the PI and PG steps.
According to Fig. 9, the LOPE’s average ground-truth reward dramatically increases at the beginning of training, similar to PPO’s learning trend in the default OpenAI Gym reward setting. Meanwhile, the value of the LOPE’s average ground-truth reward improves incrementally during policy optimization, and this value gradually approximates that of PPO with the default reward function by the end of training. This result suggests that human feedback can facilitate the agent’s learning of near-optimal behaviors that are rewarded with a high score by the default OpenAI Gym reward function.
To obtain a more intuitive presentation of the trajectories learned by the agent, we display the behaviors of agents trained using LOPE with human preferences and PPO with dense default rewards. Interestingly, LOPE can learn similar behavior patterns compared with those trained with PPO. This fact demonstrates that simple trajectory-wise state marginal matching can guide the agent in generating near-optimal behaviors, thereby avoiding the need for complicated reward engineering.
@@ VII-E Ablation Analyses
The experimental results presented in the previous section indicate that LOPE outperforms other baseline approaches on several challenging tasks. We are now interested in whether these advantages still hold when the preference setting changes. We conducted the ablation experiments in five different aspects. We first conducted another ablation study on the algorithm structure to demonstrate the effectiveness of preference guidance. We then designed an ablation experiment and compared the LOPE’s performance on preferences with different qualities. Moreover, we exchanged the PI and the PG steps and observed the experimental results. Finally, we considered using different kernel functions and combining the weakly supervised learning technique to test the performance of LOPE. These ablation studies are based on the grid world environment.
@@
LOPE with different structures. We ran LOPE without the policy improvement step (LOPE w/o PI) and LOPE without the preference guidance step (LOPE w/o PG) and obtained their results separately. The results are presented in Fig. 11LABEL:sub@fig:ablation. LOPE w/o PG cannot find the treasure and learn the optimal policy. In contrast, LOPE w/o PI can find the treasure; however, the learning efficiency of the optimal policy is low. This result implies that preference guidance can enhance the agent’s exploration efficiency, while the policy improvement step ensures the reproduction of these successful trajectories.
Preferences with different qualities. In this experiment, preference-labeled trajectories can be mislabeled with different ratios. Specifically, we experimented with synthetic noisy labels in the preference comparator (e.g., by randomly flipping a portion of pairwise feedback with ratios $0,0.1,0.2,0.3$) and obtained corresponding experimental results. The performance comparison is shown in Fig. 11b. LOPE achieves the highest performance when the mislabeling ratio is 0. Moreover, we observed that LOPE remains relatively stable up to moderate noise levels ($\approx$20%). These early results suggest that the implicit averaging effect of trajectory-level marginal matching provides a degree of robustness against noisy labels.
Exchanging the PI and PG steps. We exchanged the PI and PG steps of LOPE and compared their performance in the grid-world task with random goals. ”PI+PG” represents the LOPE version that performs the PI step first, while ”PG+PI” represents the opposite. The results are presented in Fig. 11c. The experimental result illustrates that these two methods can achieve approximately the same final performance. However, the PI+PG version learns faster than PG+PI in the early stage of the training, likely because updating the policy with stronger environmental signals (PI step) first establishes a better exploration foundation, upon which human preferences can more effectively guide further behavior..
Ablation experiment of weakly supervised learning. We combined the weakly supervised learning technique to leverage unlabeled data to assist in the policy optimization process in the initial stage of training. A cross-entropy (CE) loss is added to the policy improvement objective function because of limited supervised signal information. The experimental result is shown in Fig. 12a. This result demonstrates that the weakly supervised learning technique makes a significant contribution to improving the algorithm’s performance during the early stages of training. A possible reason is that the weakly supervised learning loss function can increase the likelihood of the preferred action, encouraging the agent to visit promising regions with higher rewards.
Figure 12: (a) The performance comparison between the original LOPE and the LOPE augmented with weakly supervised learning; (b) Ablation results of different kernel functions.
Ablation results of different kernel choices. We conducted an ablation study to assess the impact of different kernel functions used in the MMD-based preference-guided state marginal matching objective. Specifically, we evaluated four representative kernels: linear, Gaussian, Laplacian, and multiquadric. As shown in Fig. 12b, the Laplacian kernel delivers the most favorable performance, achieving both fast convergence and high final success rate. The Gaussian kernel follows closely, exhibiting similarly strong results. The multiquadric kernel leads to slower learning and plateaus at a lower success rate, indicating poor generalization to preferred trajectories. Notably, the linear kernel completely fails to guide policy learning, resulting in zero improvement throughout training. This highlights the importance of nonlinearity and localized sensitivity in kernel selection. Overall, the experiment suggests that Laplacian and Gaussian kernels are most effective for modeling trajectory-level preferences under our marginal matching framework.
@@ VIII Conclusion
This study develops a practical PbRL algorithm for hard-exploration tasks with long horizons and sparse rewards. Our key idea is to adjust the agent’s exploration by considering preferred trajectories as guidance, thereby avoiding the need to learn a reward function. LOPE involves a two-step sequential policy optimization process, including trust-region-based policy improvement and preference guidance steps. Preference guidance is reformulated as a novel trajectory-wise state marginal matching problem that minimizes a novel maximum mean discrepancy distance between the preferred trajectories and the learned policy. We further provide a theoretical analysis to characterize the performance improvement bound and evaluate the effectiveness of the proposed approach. This result implies convergence to a near-optimal policy under mild assumptions on the preference quality and KL constraints, thus offering both performance insight and theoretical justification of LOPE’s behavior. Extensive experimental results verify the superior performance of LOPE over different competitive baseline methods on several sparse-reward benchmark tasks.
In future work, we aim to enhance the robustness and generality of LOPE further. One important direction is to enhance the system’s reliability in the face of noisy, conflicting, or low-quality human preferences. While LOPE exhibits preliminary resilience under moderate synthetic noise, incorporating uncertainty-aware preference modeling (e.g., confidence-weighted aggregation or disagreement filtering) could significantly improve robustness in real-world deployments. Another promising extension is to integrate multimodal human feedback, such as natural language descriptions or voice instructions, to enrich the agent’s understanding of task intent. LOPE’s modular framework naturally accommodates these enhancements, and we consider these challenges to be important next steps toward building more practical and human-aligned preference-based reinforcement learning systems.
@@ Appendix A Proof of Lemma 1
$\underset{\tau\sim\rho_{\theta}}{\mathbb{E}}\left[\underset{\upsilon\sim\mathcal{P}}{\mathbb{E}}\left[d(\tau,\upsilon)\right]\right]$ $=\underset{\tau\sim\rho_{\theta}}{\mathbb{E}}\left[\underset{\upsilon\sim\mathcal{P}}{\mathbb{E}}\left[\underset{s_{\tau}\sim\rho_{\tau}}{\mathbb{E}}\left[\operatorname{dist}(s_{\tau},\upsilon)\right]\right]\right]$ (18)
$=\underset{\tau\sim\rho_{\theta}}{\mathbb{E}}\left[\underset{s_{\tau}\sim\rho_{\tau}}{\mathbb{E}}\left[\underset{\upsilon\sim\mathcal{P}}{\mathbb{E}}\left[\operatorname{dist}(s_{\tau},\upsilon)\right]\right]\right],$
where $d_{\theta}$ is the discounted state visitation distribution defined in Section III. In this manner, $\underset{\upsilon\sim\mathcal{P}}{\mathbb{E}}\left[\operatorname{dist}(s_{\tau},\upsilon)\right]$ in Eq. (18) can be viewed as an intrinsic reward and let $r_{g}(s,a)=\underset{\upsilon\sim\mathcal{P}}{\mathbb{E}}\left[\operatorname{dist}(s_{\tau},\upsilon)\right]$. Then, we obtain
$\underset{\begin{subarray}{l}s\sim d_{\theta}\\ a\sim\pi_{\theta}\end{subarray}}{\mathbb{E}}\left[r_{g}(s,a)\right]=\underset{\tau\sim\rho_{\theta}}{\mathbb{E}}\left[\underset{\upsilon\sim\mathcal{P}}{\mathbb{E}}\left[d(\tau,\upsilon)\right]\right].$ (19)
@@ Appendix B Proof of Policy Performance Bound
Definition 1 ($\alpha$-coupled policies [schulman2015trust, kang2018policy]).
[…]
@@
Lemma 3 (Adopted from [levin2017markov]).
@@
Lemma 4 (Adopted from [schulman2015trust, achiam2017constrained]).
@@
The proof details can be found in [schulman2015trust], and we omit them here for brevity. Given two arbitrary stochastic policies $\pi_{1}$ and $\pi_{2}$, define the expected advantage of the policy $\pi_{1}$ over the policy $\pi_{2}$ at state $s$ as $A^{\pi_{1}|\pi_{2}}(s)$:
@@
[…]
@@
A similar result can be found in [kang2018policy].
Lemma 5 (Adopted from [kang2018policy]).
@@
The proof details of Lemma 5 can be found in [kang2018policy], and we omit them here.
@@
Now, we will derive the performance improvement bound described in Lemma 2. The proof of Lemma 2 is described below.
Proof of Lemma 2.
Since $\beta=D_{\rm KL}^{\max}(\tilde{\pi},\pi_{b})=\max_{s}D_{\rm KL}(\tilde{\pi}(\cdot|s),\pi_{b}(\cdot|s))$, considering $D_{\rm TV}^{2}(p,q)\leq D_{\rm KL}(p,q)$ [pollard2000asymptopia], we assume $\alpha=D_{\rm TV}^{\max}(\tilde{\pi},\pi_{b})$ and $\alpha$ satisfies $\alpha^{2}\leq\beta$. According to Lemma 3, $\pi_{b}$, $\tilde{\pi}$ are $\alpha$-coupled. Furthermore, it is worth mentioning that the expert policy $\pi_{b}$ satisfies Assumption 1, which means
@@
[…]
@@
$\geq$ $-\sum^{\infty}_{t=0}\gamma^{t}\left(2\alpha(1-(1-\alpha)^{t})\max_{s,a}|A_{\pi_{b}}(s,a)|-\Delta\right)$ (37)
@@
where $\epsilon_{b}=\max_{s,a}|A_{b}(s,a)|$ is the maximum absolute value of $\pi_{b}$’s advantage function $A_{b}(s,a)$. Noting $\beta=D_{\rm KL}^{\max}(\tilde{\pi},\pi_{b})$ and applying $D_{\rm TV}^{2}(p,q)\leq D_{\rm KL}(p,q)$ [pollard2000asymptopia] gives
@@
TABLE I: Hyperparameter choices of LOPE for different tasks.
Hyperparameter GridWorld AntMaze HalfCheetah Hopper PointMaze-Medium PointMaze-Large
Value Function Learning Rate $0.01$ $0.001$ $0.001$ $0.001$ 1.2e-4 1.2e-4
Policy Learning Rate $0.000015$ $0.00009$ $0.00009$ $0.0004$ 1.8e-5 1.8e-5
Actor Regularization None None None None None None
Policy Optimizer Adam Adam Adam Adam Adam Adam
Value Function Optimizer Adam Adam Adam Adam Adam Adam
Max Length per Episode $240$ $750$ $500$ $500$ 250 600
Episodes per iteration 88 $30$ 2020 2020 12 12
Epochs per iteration $80$ $65$ $80$ $80$ 65 65
Discount Factor $0.99$ $0.99$ $0.99$ $0.99$ 0.99 0.99
Normalized Observations False False False False False False
Gradient Clipping False False False False False False
Initial Exploration Policy None $N(0,0.17)$ $N(0,0.15)$ $N(0,0.65)$ $N(0,0.65)$ $N(0,0.75)$
Clipped Epsilon $0.3$ $0.21$ $0.2$ $0.27$ 0.4 0.4
Finally, we describe the proof of Theorem 1.
Proof of Theorem 1.
@@
Appendix C Neural Network Architectures and Hyperparameters
C-A Neural Network Architectures
LOPE policy architecture for discrete control tasks: